# H-1117 Budapest, Pázmány Péter sétány 1/A, Hungary

### v1

@@ 4 Discussion
Figure 2: The gas density distribution at the end of the simulation is normalized with respect to the initial distribution around an Earth-mass planet for $\mathrm{St}=0.01$ in a high-metallicity disk $\epsilon=0.1$. The three different initial disk density profiles with steepness values of $p=0.5,1.0,$ and 1.5, and the three different accretion efficiencies with values of $\eta=0\%,10\%,$ and $100\%,$ are shown. The circles represent the planetary Hill sphere. The orbital motion of the planet is indicated.
Figure 3: The final gas density distribution is normalized by that in the canonical metallicity model in the vicinity of an Earth-mass planet, assuming a solid species with $\mathrm{St}=0.01$ in a disk with a metallicity of $0.03$ and $0.1$, respectively, and a density profile of $p=0.5$.
Figure 4: The difference in the azimuthally averaged radial gas torque profiles is shown for species with $\mathrm{St}=0.01$ in disks of varying metallicity. The different accretion efficiencies $\eta=0\%,10\%$, and $100\%$ are represented by the colors black, blue and red, respectively.
In this section, we examine models in which the total torque prediction deviates not only quantitatively but also qualitatively from the outcomes of numerical simulations. Such discrepancies are found for solid species with $\mathrm{St}=0.01$ and $\mathrm{St}=1$.
@@
Figure 3 shows the gas surface density around the planet for $\mathrm{St}=0.01$ in $p=0.5$ disks with high metallicity, normalized to the corresponding canonical metallicity distributions. At $\epsilon=0.03$ (top panels), the gas density is only slightly enhanced ahead of the planet. At $\epsilon=0.1$ (bottom panels), a more pronounced and symmetric enhancement develops around the $Y$-axis for non-accreting ($\eta=0\%$) and moderately accreting ($\eta=10\%$) planets. In the case of strong accretion ($\eta=100\%$), however, a notable gas accumulation emerges behind the planet. This asymmetry is attributed to the enhanced back-reaction from the solid component of the disk, which significantly perturbs the local gas dynamics. A similar pattern is observed in $p=1.0$ disks, while it is absent in the $p=1.5$ models.
We quantified the deviation in gas torque between high- and canonical metallicity disks in the planet’s vicinity via:
@@
Figure 5: The gas density distribution at the end of the simulation is normalized with respect to the initial distribution around an Earth-mass planet for $\mathrm{St}=1$ in disk assuming metallicity of $\epsilon=0.01,\penalty 10000\ 0.03,$ and $0.1$ and $p=0.5$. The three different accretion efficiencies with values of $\eta=0\%,10\%,$ and $100\%,$ are shown. The circles represent the planetary Hill sphere. The orbital motion of the planet is indicated.
Figure 6: The final gas density distribution is normalized by that in the canonical metallicity model in the vicinity of an Earth-mass planet, assuming a solid species with $\mathrm{St}=1$ in a disk with a metallicity of $0.03$ and $0.1$, respectively, and a density profile of $p=0.5$.
Figure 7: The difference in the azimuthally averaged radial gas torque profiles is shown for species with $\mathrm{St}=1$ in disks of varying metallicity. The different accretion efficiencies $\eta=0\%,10\%$, and $100\%$, are represented by the colors black, blue and red, respectively.
@@
Figure 6 shows the normalized gas surface density in the $p=0.5$ disk for high metallicities, relative to the canonical metallicity case for $\mathrm{St}=1$. At $\epsilon=0.03$ (top panels), the gas distribution exhibits only slight enhancement along the planetary spiral waves, with a modest depression in the co-orbital region trailing the planet. However, at higher metallicity ($\epsilon=0.1$, bottom panels), the trailing depletion becomes substantially deeper. Additionally, an asymmetry emerges in the spiral waves: the leading spiral wave shows a marked density enhancement, while the trailing wave exhibits a corresponding depletion. This asymmetry also contributes to the gas torque becoming positive in the high-metallicity ($\epsilon=0.1$) case. These features are characteristic signatures of the disk’s increased metallicity.
@@ 5 Caveats
Most importantly, the planet was held on a fixed orbit and thus only torques were computed. Paardekooper (2014) formalizes the distinction of static (migration-rate independent) vs dynamical (migration-rate dependent) components for low-mass planets and shows that dynamical corotation torques arise when the planet actually migrates – these can strongly modify both magnitude and sign relative to the static torque. Numerous studies have demonstrated that planetary migration can alter the distribution and accumulation of solid material in the planet’s vicinity, thereby exerting a substantial impact on the surrounding gas dynamics (see, e.g., Meru et al., 2019; Wafflard-Fernandez and Baruteau, 2020; Kanagawa et al., 2021; Kuwahara et al., 2024). For instance, in a recent work Chametla et al. (2025a) showed that a dust-depleted region trailing the planet can drive either runaway migration or oscillatory torque-driven migration.
Another limitation of our study is that it focused exclusively on planets with a mass of $1\penalty 10000\ M_{\oplus}$, for which the influence of the solid component on the total torque is most pronounced (Regály et al., 2025). Exploring a broader range of planetary masses would provide insight into how solid back-reaction in high-metallicity disks scales with planetary mass. Furthermore, modeling particles with large Stokes numbers ($\mathrm{St}>1$) may require a Lagrangian treatment, as demonstrated by Chrenko et al. (2024). Our simulations were performed in two dimensions; however, fully three-dimensional hydrodynamic models may uncover additional phenomena. For example, Chametla et al. (2025a) demonstrated that the structure of the dust-depleted region (dust-void) is significantly altered in three dimensions. Finally, we employed a constant Stokes number approximation, which assumption may be inaccurate in regions with strong gas density perturbations, where the local gas properties – thus the effective Stokes number for a given particle size – can vary significantly. Future work should aim to address these simplifications to provide a more realistic and predictive framework for the study of low-mass planetary migration.
@@ 6 Conclusions
We investigated the effect of disk metallicity, modeled by changing the solid-to-gas mass ratio of the disk ($\epsilon$), on the torque exerted on a $1\penalty 10000\ M_{\oplus}$ planet by the disk’s solid and gaseous components. The planet was maintained on a fixed circular orbit, and consequently, only static torques were computed. These may differ from the torques that would arise in a fully dynamical (migrating) configuration. We investigated higher disk metallicities in the cases of $\epsilon=0.03$ and $\epsilon=0.1$ and compared them to the canonical case of $\epsilon=0.01$. Based on our earlier models (Regály et al., 2025), the torque components were predicted by considering both the torque exerted by the solid component and its back-reaction on the gas dynamics. To verify these predictions, we conducted two-dimensional numerical hydrodynamic simulations of planet-disk interactions in high-metallicity disks. In these simulations, we took into account the back-reaction of the solid disk component on the gas dynamics. Our main findings are as follows:
@@
    The torque contribution from the solid component of the disk is qualitatively well reproduced by theoretical predictions. In contrast, the gas torque often exhibits substantial quantitative discrepancies – ranging from 50% to 100% – and, in some cases, even qualitative differences such as a reversal in sign. Consequently, the total torque experienced by a $1\penalty 10000\ M_{\oplus}$ planet, arising from both gas and solid components, can only be reliably predicted in high-metallicity disks for particles with $\mathrm{St}\geq 3$, irrespective of the planetary accretion efficiency.
@@
    For species with $\mathrm{St}=1$, the predictions yield a very strong negative torque. However, the numerical simulations result in a total torque magnitude that is an order of magnitude smaller ($p=0.5$) or half the magnitude ($p=1.0$). This phenomenon also occurs for solid species with $\mathrm{St}=2$ in disks with a steep surface density profile ($p=1.5$).
@@
    For coupled solid species with $\mathrm{St}=0.01$, numerical simulations result in negative total torque when planetary accretion is significant ($10\%$ or $100\%$) and the surface density slope is shallow ($p=0.5$ or $1.0$). However, the predictions suggest strong positive total torque.
Previously we showed that the gas torque is significantly altered due to the back-reaction of solid species (Regály et al., 2025). Since the effect of back-reaction of solid species is enhanced in high-metallicity disks, the resulting change in gas torque significantly alters the total torque felt by the planet. Thus care must be taken when extrapolating the total torque from canonical metallicity ($\epsilon=0.01$) disk simulations. The simplified assumptions that the torque is simply multiplied by the elevated solid-to-gas mass ratio is not valid for all Stokes number investigated ($0.01\leq\mathrm{St}\leq 10$). In light of our findings, migration studies of low-mass planets ($0.1\penalty 10000\ M_{\oplus}\leq M_{\mathrm{p}}\leq 10\penalty 10000\ M_{\oplus}$) requires numerical simulations using proper solid content in the disk instead of predicting the torque from canonical metallicity disk models.
@@ Acknowledgements.
We thank the anonymous referee, whose helpful comments significantly helped to improve the quality of this paper.

### v2

@@ 4 Discussion
Figure 2: Gas density distribution at the end of the simulation normalized with respect to the initial distribution around an Earth-mass planet for $\mathrm{St}=0.01$ in a high-metallicity disk with $\epsilon=0.1$. The three different initial disk density profiles with steepness values of $p=0.5,1.0,$ and 1.5, and the three different accretion efficiencies with values of $\eta=0\%,10\%,$ and $100\%,$ are shown. The circles represent the planetary Hill sphere. The orbital motion of the planet is indicated.
Figure 3: Final gas density distribution normalized by that in the canonical-metallicity model in the vicinity of an Earth-mass planet, assuming a solid species with $\mathrm{St}=0.01$ in a disk with a metallicity of $0.03$ (top) and $0.1$ (bottom) and a density profile of $p=0.5$.
Figure 4: Difference in the azimuthally averaged radial gas torque profiles for species with $\mathrm{St}=0.01$ in disks of varying metallicities. The different accretion efficiencies $\eta=0\%,10\%$, and $100\%$ are shown in black, blue, and red, respectively.
In this section we examine models in which the total torque prediction deviates not only quantitatively but also qualitatively from the outcomes of numerical simulations. Such discrepancies are found for solid species with $\mathrm{St}=0.01$ and $\mathrm{St}=1$.
@@
Figure 3 shows the gas surface density around the planet for $\mathrm{St}=0.01$ in $p=0.5$ disks with high metallicity, normalized to the corresponding canonical-metallicity distributions. At $\epsilon=0.03$ (top panels), the gas density is only slightly enhanced ahead of the planet. At $\epsilon=0.1$ (bottom panels), a more pronounced and symmetric enhancement develops around the $Y$-axis for non-accreting ($\eta=0\%$) and moderately accreting ($\eta=10\%$) planets. In the case of strong accretion ($\eta=100\%$), however, a notable gas accumulation emerges behind the planet. This asymmetry is attributed to the enhanced back-reaction from the solid component of the disk, which significantly perturbs the local gas dynamics. A similar pattern is observed in $p=1.0$ disks, while it is absent in the $p=1.5$ models.
We quantified the deviation in gas torque between high- and canonical-metallicity disks in the planet’s vicinity via
@@
Figure 5: Gas density distribution at the end of the simulation normalized with respect to the initial distribution around an Earth-mass planet for $\mathrm{St}=1$ in disk assuming metallicities of $\epsilon=0.01,\penalty 10000\ 0.03,$ and $0.1$ and $p=0.5$. The three different accretion efficiencies, $\eta=0\%,10\%,$ and $100\%,$ are shown. The circles represent the planetary Hill sphere. The orbital motion of the planet is indicated.
Figure 6: Final gas density distribution normalized by that in the canonical-metallicity model in the vicinity of an Earth-mass planet, assuming a solid species with $\mathrm{St}=1$ in a disk with a metallicity of $0.03$ (top) and $0.1$ (bottom) and a density profile of $p=0.5$.
Figure 7: Difference in the azimuthally averaged radial gas torque profiles for species with $\mathrm{St}=1$ in disks of varying metallicities. The different accretion efficiencies $\eta=0\%,10\%$, and $100\%$ are shown in black, blue, and red, respectively.
@@
Figure 6 shows the normalized gas surface density in the $p=0.5$ disk for high metallicities, relative to the canonical-metallicity case for $\mathrm{St}=1$. At $\epsilon=0.03$ (top panels), the gas distribution exhibits only slight enhancement along the planetary spiral waves, with a modest depression in the co-orbital region trailing the planet. However, at higher metallicity ($\epsilon=0.1$, bottom panels), the trailing depletion becomes substantially deeper. Additionally, an asymmetry emerges in the spiral waves: the leading spiral wave shows a marked density enhancement, while the trailing wave exhibits a corresponding depletion. This asymmetry also contributes to the gas torque becoming positive in the high-metallicity ($\epsilon=0.1$) case. These features are characteristic signatures of the disk’s increased metallicity.
@@ 5 Caveats
Most importantly, the planet was held on a fixed orbit and thus only torques were computed. Paardekooper (2014) formalizes the distinction of static (migration-rate-independent) versus dynamical (migration-rate-dependent) components for low-mass planets and shows that dynamical corotation torques arise when the planet actually migrates – these can strongly modify both the magnitude and sign relative to the static torque. Numerous studies have demonstrated that planetary migration can alter the distribution and accumulation of solid material in the planet’s vicinity, thereby exerting a substantial impact on the surrounding gas dynamics (see, e.g., Meru et al., 2019; Wafflard-Fernandez and Baruteau, 2020; Kanagawa et al., 2021; Kuwahara et al., 2024). For instance, in a recent work Chametla et al. (2025a) showed that a dust-depleted region trailing the planet can drive either runaway migration or oscillatory torque-driven migration.
Another limitation of our study is that it focused exclusively on planets with a mass of $1\penalty 10000\ M_{\oplus}$, for which the influence of the solid component on the total torque is most pronounced (Regály et al., 2025). Exploring a broader range of planetary masses would provide insight into how solid back-reaction in high-metallicity disks scales with planetary mass. Furthermore, modeling particles with large Stokes numbers ($\mathrm{St}>1$) may require a Lagrangian treatment, as demonstrated by Chrenko et al. (2024). Our simulations were performed in two dimensions; however, fully three-dimensional hydrodynamic models may uncover additional phenomena. For example, Chametla et al. (2025a) demonstrated that the structure of the dust-depleted region (dust-void) is significantly altered in three dimensions. Finally, we employed a constant Stokes number approximation; this assumption may be inaccurate in regions with strong gas density perturbations, where the local gas properties – and thus the effective Stokes number for a given particle size – can vary significantly. Future work should aim to address these simplifications to provide a more realistic and predictive framework for the study of low-mass planetary migration.
@@ 6 Conclusions
We investigated the effect of disk metallicity, modeled by changing the solid-to-gas mass ratio of the disk ($\epsilon$), on the torque exerted on a $1\penalty 10000\ M_{\oplus}$ planet by the disk’s solid and gaseous components. The planet was maintained on a fixed circular orbit, and consequently, only static torques were computed. These may differ from the torques that would arise in a fully dynamical (migrating) configuration. We investigated higher disk metallicities in the cases of $\epsilon=0.03$ and $\epsilon=0.1$ and compared them to the canonical case of $\epsilon=0.01$. Based on our earlier models (Regály et al., 2025), the torque components were predicted by considering both the torque exerted by the solid component and its back-reaction on the gas dynamics. To verify these predictions, we conducted two-dimensional numerical hydrodynamic simulations of planet–disk interactions in high-metallicity disks. In these simulations, we took the back-reaction of the solid disk component on the gas dynamics into account. Our main findings are as follows:
@@
    The torque contribution from the solid component of the disk is qualitatively well reproduced by theoretical predictions. In contrast, the gas torque often exhibits substantial quantitative discrepancies – ranging from 50% to 100% – and, in some cases, even qualitative differences, such as a reversal in sign. Consequently, the total torque experienced by a $1\penalty 10000\ M_{\oplus}$ planet, arising from both gas and solid components, can only be reliably predicted in high-metallicity disks for particles with $\mathrm{St}\geq 3$, irrespective of the planetary accretion efficiency.
@@
    For species with $\mathrm{St}=1$, the predictions yield a very strong negative torque. However, the numerical simulations yield a total torque magnitude that is an order of magnitude smaller ($p=0.5$) or half the magnitude ($p=1.0$). This phenomenon also occurs for solid species with $\mathrm{St}=2$ in disks with a steep surface density profile ($p=1.5$).
@@
    For coupled solid species with $\mathrm{St}=0.01$, numerical simulations result in a negative total torque when planetary accretion is significant ($10\%$ or $100\%$) and the surface density slope is shallow ($p=0.5$ or $1.0$). However, the predictions suggest a strong positive total torque.
Previously we showed that the gas torque is significantly altered due to the back-reaction of solid species (Regály et al., 2025). Since the effect of the back-reaction of solid species is enhanced in high-metallicity disks, the resulting change in the gas torque significantly alters the total torque felt by the planet. Thus, care must be taken when extrapolating the total torque from canonical-metallicity ($\epsilon=0.01$) disk simulations. The simplified assumptions that the torque is simply multiplied by the elevated solid-to-gas mass ratio is not valid for all Stokes number investigated ($0.01\leq\mathrm{St}\leq 10$). In light of our findings, migration studies of low-mass planets ($0.1\penalty 10000\ M_{\oplus}\leq M_{\mathrm{p}}\leq 10\penalty 10000\ M_{\oplus}$) require numerical simulations that use proper solid content in the disk instead of predicting the torque from canonical-metallicity disk models.
@@ Acknowledgements.
We thank the anonymous referee, whose helpful comments significantly helped improve the quality of this paper.